# SketchGuard: Scaling Byzantine-Robust Decentralized Federated Learning via Sketch-Based Screening

### v2

@@ B.1. Starting from Smoothness
(26) $\quad+2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$
B.2. Rearranging for Gradient Norm
@@
(27) $\quad+2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$
B.3. Telescoping
@@
(28) $\quad+T(2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha))$
B.4. Averaging and Using Lower Bound
@@
(29) $\quad+4L\eta\delta^{2}+\frac{4\gamma_{\text{eff}}\psi\rho(1-\alpha)}{\eta}$
This establishes Theorem 2. The convergence rate matches the optimal rate for non-convex optimization, with the compression error appearing only through $\gamma_{\text{eff}}$.
Appendix C Supplementary Experimental Details
This appendix provides comprehensive details about the experimental setup, including model architectures, hyperparameter configurations, network topologies, and scalability experiment specifications used throughout our study.
C.1. Detailed Model Architectures
We provide complete architectural specifications for the two datasets used in our experiments: FEMNIST and CelebA. Both architectures follow standard designs adapted to their respective input dimensions and task requirements.
C.1.1. FEMNIST Architecture
@@
Table 4. FEMNIST Model Architecture Details
@@ C.1.2. CelebA Architecture
Table 5. CelebA Model Architecture Details
@@
C.2. Complete Experimental Configuration
@@
C.2.1. Network Topology Specifications
@@
Table 6. Network Topology Configurations
@@
C.2.2. Hyperparameter Settings
Table 7 presents the complete set of hyperparameters used in our experiments, organized by category. Training configuration parameters were selected based on standard federated learning practices. Defense mechanism parameters ($\gamma$, $\kappa$, $\alpha$) were tuned to balance robustness and model utility. For SketchGuard, we set the sketch size $k=1000$ for FEMNIST and $k=350$ for CelebA to get roughly the same factor of compression for both datasets.
Table 7. Complete Hyperparameter Configuration
@@ C.2.2. Hyperparameter Settings
Learning rate 0.01 (FEMNIST), 0.001 (CelebA)
Maximum samples per client 10,000 (FEMNIST), 4,500 (CelebA)
Random seed 987654321, 39573295, 32599368
@@
Sketch size ($k$) 1,000 (FEMNIST), 350 (CelebA)
@@
C.3. Scalability Experiments
@@
C.3.1. Network Size Scaling
To evaluate the computational scalability of defense mechanisms with respect to network size, we conducted experiments on k-regular graphs with varying numbers of participants. Table 8 shows the configurations used, with network sizes ranging from 20 to 300 nodes. All networks maintain 50% Byzantine nodes to test defense robustness under high attack scenarios. The node degree was selected to ensure connectivity while maintaining realistic peer-to-peer network constraints. These experiments used shorter training runs (3 rounds with 1 local epoch each) to focus on measuring computational overhead rather than convergence behavior.
Table 8. Network Size Scaling Configurations
Node Degree Network Size Attack Percentage
@@ C.3.1. Network Size Scaling
C.3.2. Model Scaling Variants
@@
Table 9. Model Scaling Variants
@@
Standard Baseline FEMNIST architecture 6,603,710

### v3

@@ B.1 Starting from Smoothness
$\quad+2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$ (26)
B.2 Rearranging for Gradient Norm
@@
$\quad+2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$ (27)
B.3 Telescoping
@@
$\quad+T(2L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha))$ (28)
B.4 Averaging and Using Lower Bound
@@
$\quad+4L\eta\delta^{2}+\frac{4\gamma_{\text{eff}}\psi\rho(1-\alpha)}{\eta}$ (29)
This establishes Theorem 4.8. The convergence rate matches the optimal rate for non-convex optimization, with the compression error appearing only through $\gamma_{\text{eff}}$.
Appendix C Comparison with Existing Methods
Table 3 compares SketchGuard with existing Byzantine-robust DFL algorithms across key properties. SketchGuard is the only method that combines convergence guarantees in both convex and non-convex settings, operates without requiring knowledge of the compromised node ratio or complete graph assumptions, and achieves high scalability through sketch-based compression.
Table 3: Comparison of Byzantine-robust decentralized aggregation algorithms. ”Convex”/”Non-convex” indicate convergence guarantees; ”No $c_{i}$” means no need to know compromised node ratio; ”No Complete” means no complete-graph assumption.
Algorithm Convex Non-convex No $c_{i}$ No Complete Scalable
UBAR (Guo et al., 2022) Med.
LEARN (El-Mhamdi et al., 2021) Low
SCCLIP (He et al., 2022) Med.
BALANCE (Fang et al., 2024) Med.
WFAgg (Cajaraville-Aboy et al., 2025) Med.
SketchGuard High
Appendix D Supplementary Experimental Details
This appendix provides comprehensive details about the experimental setup, including model architectures, hyperparameter configurations, network topologies, attack configurations, and scalability experiment specifications used throughout our study.
D.1 Detailed Model Architectures
We provide complete architectural specifications for the three datasets used in our experiments: FEMNIST, CelebA, and Sent140. The image-based models (FEMNIST and CelebA) follow standard convolutional designs, while Sent140 uses an LSTM-based architecture for sequential text data.
D.1.1 FEMNIST Architecture
@@
Table 4: FEMNIST Model Architecture Details
@@ D.1.2 CelebA Architecture
Table 5: CelebA Model Architecture Details
@@
D.1.3 Sent140 Architecture
The Sent140 model uses a stacked LSTM architecture following the LEAF reference implementation. The model processes variable-length text sequences through an embedding layer, a 2-layer LSTM, and fully connected output layers. Table 6 provides the complete specification.
Table 6: Sent140 Model Architecture Details
Layer Type Configuration Output Size
Input seq_len
Embedding Embedding vocab=10000, dim=100 $\text{seq\_len}\times 100$
LSTM Stacked LSTM 2 layers, hidden=100 $100$
Dropout Dropout $p=0.1$ $100$
FC1 Linear $100\rightarrow 128$ $128$
FC2 Linear $128\rightarrow 2$ 22
Total $\sim$1.2M
The embedding layer maps vocabulary indices (size 10,000) to 100-dimensional dense vectors. The stacked 2-layer LSTM with 100 hidden units processes the embedded sequence with dropout for regularization. The final hidden state passes through two fully connected layers for binary sentiment classification.
D.2 Evaluation Metrics
We evaluate defense mechanisms using four complementary metrics that capture robustness, backdoor resilience, computational efficiency, and communication overhead.
Test Error Rate (TER). Defined as 1test accuracy1-$\text${test accuracy}, TER measures performance degradation averaged across honest clients. Lower values indicate better Byzantine resilience against model poisoning attacks. TER is computed on each honest client’s local test set and averaged across all honest clients in the network.
Attack Success Rate (ASR). For backdoor attacks, ASR measures the fraction of triggered test samples classified as the attacker’s target label on honest client models. Lower ASR indicates better resilience to backdoor injection. The random-chance baseline is ASR $\approx 1/C$ where $C$ is the number of classes: 1.6% for FEMNIST (62 classes) and 50% for CelebA and Sent140 (binary classification). ASR is computed by applying the backdoor trigger to clean test samples and measuring the fraction misclassified to the target label.
Per-Round Computation Time. Wall-clock time for neighbor screening and aggregation phases per client per round, excluding local training which remains constant across methods. This metric captures the computational overhead introduced by Byzantine-robust aggregation mechanisms. All timing measurements are recorded on identical hardware (CPU-only for reproducibility) and averaged across all clients and rounds.
Communication Overhead. Total parameters transmitted per client per round, including both sketches (for SketchGuard) and full model exchanges. We report both absolute parameter counts and reduction percentages relative to full-precision baselines that transmit complete models to all neighbors.
D.3 Complete Experimental Configuration
@@
D.3.1 Network Topology Specifications
@@
Table 7: Network Topology Configurations
@@
D.3.2 Hyperparameter Settings
Table 8 presents the complete set of hyperparameters used in our experiments, organized by category. Training configuration parameters were selected based on standard federated learning practices. Defense mechanism parameters ($\gamma$, $\kappa$, $\alpha$) were tuned to balance robustness and model utility. For SketchGuard, we set dataset-specific sketch sizes to achieve comparable compression ratios across different model dimensions.
Table 8: Complete Hyperparameter Configuration
@@ D.3.2 Hyperparameter Settings
Learning rate 0.01 (FEMNIST, Sent140), 0.001 (CelebA)
Maximum samples per client 10,000 (FEMNIST), 4,500 (CelebA, Sent140)
Random seeds 987654321, 39573295, 32599368
@@
Sketch size ($k$) 1,000 (FEMNIST), 350 (CelebA), 180 (Sent140)
@@
D.3.3 Backdoor Attack Configuration
Table 9 details the backdoor attack configurations used for each dataset. For image datasets (FEMNIST and CelebA), the backdoor trigger consists of a white square pattern inserted in the bottom-right corner of the image. For Sent140, the trigger consists of rare vocabulary token indices appended to the end of the text sequence.
Table 9: Backdoor Attack Configuration by Dataset
Dataset Target Label Trigger Size Trigger Type
FEMNIST 0 $4\times 4$ pixels White square (bottom-right)
CelebA 0 (not smiling) $8\times 8$ pixels White square (bottom-right)
Sent140 1 (positive) 3 tokens Rare word indices appended
For Sent140, the trigger tokens are selected from high vocabulary indices (9997, 9998, 9999) that rarely appear in natural text, ensuring the trigger pattern is distinctive and unlikely to occur in benign samples.
D.4 Detailed Attack Descriptions
We evaluate Byzantine robustness using four representative attack strategies that span different threat models and sophistication levels.
D.4.1 Directed Deviation Attack
Directed deviation attacks (Fang et al., 2020) employ an optimization-based approach that crafts malicious updates by solving for parameters that maximize deviation in the direction opposite to honest gradient descent. This adaptive attack strategy represents a strong adversary that actively attempts to subvert Byzantine-robust aggregation mechanisms by carefully positioning malicious updates to evade detection while maximizing impact. The attack computes the mean of honest updates and crafts malicious updates as:
$\hat{\mathbf{w}}_{j}=\bar{\mathbf{w}}-\lambda\cdot\text{sign}(\bar{\mathbf{w}}-\mathbf{w}^{t})$ (30)
where $\bar{\mathbf{w}}$ is the mean honest update, $\mathbf{w}^{t}$ is the current model, and $\lambda$ controls attack intensity. This attack has demonstrated effectiveness against common Byzantine-robust aggregation rules including Krum and trimmed mean.
D.4.2 Gaussian Attack
Gaussian attacks (Blanchard et al., 2017) inject random noise $\hat{\mathbf{w}}_{j}=\mathcal{N}(0,200\mathbf{I})$, representing less sophisticated but realistic adversaries that disrupt convergence through stochastic perturbations. While simpler than directed deviation, these attacks model practical scenarios where attackers lack complete knowledge of the aggregation mechanism or resources for optimization-based attacks. The high variance ($\sigma^{2}=200$) ensures significant disruption to model convergence.
D.4.3 Krum Attack
Krum attacks (Fang et al., 2020) craft malicious models that cause Krum to select the attacker’s chosen model. The attack works by:
  1. 1.
    Computing the mean of honest models as a reference
  2. 2.
    Creating a primary malicious model in the opposite direction of honest gradient descent
  3. 3.
    Creating supporting malicious models very close to the primary to minimize the Krum score of the primary model
The attack exploits Krum’s reliance on pairwise distances: if attackers control enough nodes, they can cluster their models tightly while being far from honest models, causing Krum to select the malicious model as it appears most “central.”
D.4.4 Backdoor Attack
Backdoor attacks (Bagdasaryan et al., 2020) represent a distinct threat model where adversaries aim to embed hidden functionality rather than degrade overall performance. Malicious clients augment their training data with trigger patterns mapped to an attacker-chosen target label, then scale their model updates by a factor proportional to the total number of clients to amplify the backdoor’s persistence during aggregation. Unlike other attacks that maximize test error, backdoor attacks preserve main task accuracy while causing misclassification only on triggered inputs, making them particularly stealthy. Dataset-specific trigger configurations are detailed in Table 9.
D.5 Scalability Experiments
@@
D.5.1 Network Size Scaling
To evaluate the computational scalability of defense mechanisms with respect to network size, we conducted experiments on k-regular graphs with varying numbers of participants. Table 10 shows the configurations used, with network sizes ranging from 20 to 300 nodes. All networks maintain 50% Byzantine nodes to test defense robustness under high attack scenarios. The node degree was selected to ensure connectivity while maintaining realistic peer-to-peer network constraints. These experiments used shorter training runs (3 rounds with 1 local epoch each) to focus on measuring computational overhead rather than convergence behavior. All timing measurements were performed on CPU to ensure consistent and reproducible results across different hardware configurations.
Table 10: Network Size Scaling Configurations
Node Degree ($k$) Network Size Attack Percentage
@@ D.5.1 Network Size Scaling
D.5.2 Model Scaling Variants
@@
Table 11: Model Scaling Variants
@@
Baseline Baseline FEMNIST architecture 6,603,710
@@
Model dimensionality scaling experiments use a fixed k-regular graph with node degree 154 (155 total nodes) to isolate the effect of model size from network topology effects.